# Simple Lines, Big Ideas: Towards Interpretable Assessment of Human Creativity from Drawings

### v1

@@ Abstract
Assessing human creativity through visual outputs, such as drawings, plays a critical role in fields including psychology, education, and cognitive science. However, current assessment practices still rely heavily on expert-based subjective scoring, which is both labor-intensive and inherently subjective. In this paper, we propose a data-driven framework for automatic and interpretable creativity assessment from drawings. Motivated by the cognitive understanding that creativity can emerge from both what is drawn (content) and how it is drawn (style), we reinterpret the creativity score as a function of these two complementary dimensions. Specifically, we first augment an existing creativity-labeled dataset with additional annotations targeting content categories. Based on the enriched dataset, we further propose a multi-modal, multi-task learning framework that simultaneously predicts creativity scores, categorizes content types, and extracts stylistic features. In particular, we introduce a conditional learning mechanism that enables the model to adapt its visual feature extraction by dynamically tuning it to creativity-relevant signals conditioned on the drawing’s stylistic and semantic cues. Experimental results demonstrate that our model achieves state-of-the-art performance compared to existing regression-based approaches and offers interpretable visualizations that align well with human judgments. The code and annotations will be made publicly available at https://github.com/WonderOfU9/CSCA_PRCV_2025
@@ 1 Introduction
Despite these advances, existing automated models have notable limitations. Many rely on relatively shallow architectures or simple classification schemes [zhang_painting, cropley_figural], making it hard to capture the full nuance of creative expression. In particular, black-box CNN scores are not readily interpretable [panfilova_resnet, chen_bias], and most models are trained on one specific task or dataset, raising concerns about generalization [patterson2024audra]. Models may inadvertently learn spurious cues (e.g., ink coverage) that do not reflect true creativity, leading to poor transfer across tasks [elgammal_art]. In summary, we observe three key gaps: shallow modeling, lack of interpretability, and limited cross-task robustness [gan_generalization].
To overcome these challenges, we adopt a “content + style” perspective  [Pencils_to_Pixels]. This view posits that creativity in a drawing arises from both what is depicted (content) and how it is depicted (style). For example, two sketches of a forest may contain the same content (trees and animals), but a richly textured or unusual style can make one appear more creative than the other. These observations align with findings in computational creativity and vision (e.g., neural style transfer separates content and style [gatys_styletransfer]). Therefore, instead of treating creativity as a monolithic score, we explicitly factor our model to capture content semantics (objects) and stylistic features (line complexity).
Building on this idea, our paper proposes a data-driven, interpretable framework for creativity assessment that jointly models content and style. Specifically, we design a multi-modal, multi-task neural architecture. One head of the network predicts the overall creativity score, another predicts the drawing’s content category (e.g., “plant,” “animal”, “object”), and another extracts quantifiable style descriptors (line complexity). The use of multi-task learning allows shared representations to benefit all predictions, while each task provides an interpretability signal. Crucially, we incorporate a conditional learning mechanism that dynamically modulates the image embedding based on identified content and style cues. Our main contributions are:
@@
  • We introduce a novel interpretable model for visual creativity assessment that decomposes drawings into content and style components.
@@ 3.2 Experimental results
Audra 2024 0.80 0.79
AGIQA 2025 0.79 0.80
@@
CLIPIQA+[wang2023exploring]
@@ 4 Conclusion
4.0.1 Ackowledgements

### v2

@@ Abstract
Assessing human creativity through visual outputs, such as drawings, plays a critical role in fields including psychology, education, and cognitive science. However, current assessment practices still rely heavily on expert-based subjective scoring, which is both labor-intensive and inherently subjective. In this paper, we propose a data-driven framework for automatic and interpretable creativity assessment from drawings. Motivated by the cognitive evidence proposed in [Pencils_to_Pixels] that creativity can emerge from both what is drawn (content) and how it is drawn (style), we reinterpret the creativity score as a function of these two complementary dimensions. Specifically, we first augment an existing creativity-labeled dataset with additional annotations targeting content categories. Based on the enriched dataset, we further propose a conditional model predicting content, style, and ratings simultaneously. In particular, the conditional learning mechanism that enables the model to adapt its visual feature extraction by dynamically tuning it to creativity-relevant signals conditioned on the drawing’s stylistic and semantic cues. Experimental results demonstrate that our model achieves state-of-the-art performance compared to existing regression-based approaches and offers interpretable visualizations that align well with human judgments. The code and annotations will be made publicly available at https://github.com/WonderOfU9/CSCA_PRCV_2025
@@ 1 Introduction
Despite these advances, existing automated models have notable limitations. Many rely on relatively shallow architectures or simple classification schemes [zhang_painting, cropley_figural], making it hard to capture the full nuance of creative expression. In particular, black-box CNN scores are not readily interpretable [panfilova_resnet, chen_bias], and most models are trained on one specific task or dataset, raising concerns about generalization [patterson2024audra]. Models may inadvertently learn spurious cues (e.g., ink coverage) that do not reflect true creativity, leading to poor transfer across tasks [elgammal_art]. This can also be verified by [Pencils_to_Pixels], which shows that automated tools overfit on ink density. In summary, we observe three key gaps: shallow modeling, lack of interpretability, and limited cross-task robustness [gan_generalization].
Nath et al. [Pencils_to_Pixels] present a key content–style framework for assessing creativity in drawings. They measure four stylistic attributes—ink density, fraction of ink within the stimulus shape, number of connected components, and number of lines extracted via skeletonization and Hough transforms. For content, they use CLIP image embeddings to capture semantic similarity to the base stimulus and generate short GPT-4o captions. These interpretable content and style descriptors are then combined in a regression model to predict human creativity ratings.
Building on [Pencils_to_Pixels], we design a multi-task learning architecture. One head of the network predicts the overall creativity score, another predicts the drawing’s content category (e.g., “plant,” “animal”, “object”), and another extracts quantifiable style descriptors (line complexity). The use of multi-task learning allows shared representations to benefit all predictions, while each task provides an interpretability signal. Crucially, we incorporate a conditional learning mechanism that dynamically modulates the image embedding based on identified content and style cues. Our main contributions are:
@@
  • We introduce a unified, multi-task model for visual creativity assessment that predicts the content and style components of drawings simultaneously.
@@ 3.2 Experimental results
Audra [patterson2024audra] 2024 0.80 0.79
AGIQA [tang2025clip] 2025 0.79 0.80
Pencils to Pixels [Pencils_to_Pixels] 2025 0.79 -
@@
CLIPIQA+[patterson2024audra]
@@ 4 Conclusion
4.0.1 Revision Note
This version includes expanded related work and clarifications acknowledging the contributions of Nath et al. [Pencils_to_Pixels]. The PRCV 2025 conference version was prepared earlier and does not contain these updates; the arXiv version should be considered the most complete and up-to-date reference.
4.0.2 Acknowledgments